Consider a cone with a diameter of 12 and a height of 8. The volume of an inscribed sphere in the cone is _____________
Answer by anand429(138)

http://www.algebra.com/cgi-bin/display-illustration.mpl?tutor=anand429&illustration_name=2.jpg
Hints: Tangents from an external point to a circle/sphere are equal
Slant height of cone = sqrt(h^2+r^2) = 10
Rest. see. the figure in above link.
Volume = (4/3)pi(r^3)